# Challenging Gradient Boosted Decision Trees with Tabular Transformers for Fraud Detection at Booking.com

### v1

@@ Abstract.
Transformer-based neural networks, empowered by Self-Supervised Learning (SSL), have demonstrated unprecedented performance across various domains. However, related literature suggests that tabular Transformers may struggle to outperform classical Machine Learning algorithms, such as Gradient Boosted Decision Trees (GBDT). In this paper, we aim to challenge GBDTs with tabular Transformers on a typical task faced in e-commerce, namely fraud detection. Our study is additionally motivated by the problem of selection bias, often occurring in real-life fraud detection systems. It is caused by the production system affecting which subset of traffic becomes labeled. This issue is typically addressed by sampling randomly a small part of the whole production data, referred to as a Control Group. This subset follows a target distribution of production data and therefore is usually preferred for training classification models with standard ML algorithms. Our methodology leverages the capabilities of Transformers to learn transferable representations using all available data by means of SSL, giving it an advantage over classical methods. Furthermore, we conduct large-scale experiments, pre-training tabular Transformers on vast amounts of data instances and fine-tuning them on smaller target datasets. The proposed approach outperforms heavily tuned GBDTs by a considerable margin of the Average Precision (AP) score. Pre-trained models show more consistent performance than the ones trained from scratch when fine-tuning data is limited. Moreover, they require noticeably less labeled data for reaching performance comparable to their GBDT competitor that utilizes the whole dataset.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; Under Review ; isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
  • We investigate how pre-training affects the performance of FT-Transformer and its GBDT competitor for different amounts of labeled data. We demonstrate that SSL pre-training consistently outperforms training a supervised model from scratch.
@@ 2.1. Deep Learning for Tabular Data
Recently, multiple architectures extending Transformer self-attention to tabular data have been proposed aiming to challenge and eventually overcome Gradient Boosting-based methods. In 2019, Song et al. (Song et al., 2019) introduced the AutoInt architecture converting raw tabular features to embeddings that are passed through custom self-attention layers. TabTransformer (Huang et al., 2020), introduced in 2020, exploits conventional multi-head self-attention layers (Vaswani et al., 2017) for categorical feature embeddings, whereas numerical features are processed via a separate MLP branch of the network. Eventually, the FT-Transformer architecture (Gorishniy et al., 2021) tokenizes each numerical and categorical feature, appends a class token, and applies the classical Transformer self-attention layers to the tokens. According to the original work, it outperforms the previously suggested Deep Learning models and even outperforms GBDTs on certain tasks.
@@ 4.1. Datasets
The biased (no-CG) dataset contains all instances from the traffic for which labels are available, except the ones sampled for the Control Group. The bias in this dataset is caused by a specific selection procedure, which introduces a shift in both features and labels, meaning that this dataset does not represent production data. The total number of instances in this dataset has the order of tens of millions [^8] Because of confidentiality we cannot share more precise information..
The unbiased (CG) dataset, obtained using the Control Group strategy, contains a smaller amount of instances randomly sampled from the user traffic with reconstructed labels. The number of instances in this dataset is between one and two orders of magnitude less than the no-CG dataset size8.
While the distribution of labels in the CG and no-CG datasets varies, the proportion of fraudulent examples in both datasets is very low8.
@@ 5.1. Fraud Detection
Approach Backbone Pre-training Training Test AP
Sup. LightGBM - CG 0.471 $\pm$ 0.003
Sup. LightGBM - CG + no-CG 0.441 $\pm$ 0.013
Sup. + SR LightGBM All data CG + no-CG 0.442 $\pm$ 0.011
Sup. FT-base - CG 0.463 $\pm$ 0.013
DML FT-base CG + no-CG (+) CG 0.469 $\pm$ 0.017
SSL FT-base CG + no-CG (+) CG 0.469 $\pm$ 0.013
Sup. FT-large - CG 0.454 $\pm$ 0.009
DML FT-large CG + no-CG CG 0.488 $\pm$ 0.008
SSL FT-large CG + no-CG CG 0.491 $\pm$ 0.012
@@
5.2. Smaller Unbiased Datasets
The size of the Control Group is an important hyperparameter that should be considered when implementing this strategy. A larger CG is likely to increase the amount of rare fraudulent instances in training data, which improves model performance. However, it also incurs extra costs and risks. Thus, this setting should be adjusted carefully, by taking into account both business requirements and model capabilities. To investigate this, we compared LightGBM and the large FT-Transformer, both trained from scratch and fine-tuned after SSL pre-training, in the scenarios with smaller CG sizes.
In our experiment, we randomly sampled subsets of CG with gradually increasing sizes, from 5% to 100% of CG. For each percentage, we did it 10 times with different seeds and then trained the models for each obtained subset. Eventually, we computed model performance on the same CG test set as in the previous section. The results are presented in Figure 3 as mean AP scores (solid lines) and their standard deviations (shaded areas) based on 10 runs.
Initially, when the amounts of annotated data are extremely low (5-10% of CG), all models achieve poor scores, but LightGBM demonstrates better performance. Nevertheless, this gap is then minimized between 20 and 40% of CG, where LightGBM and the pre-trained FT-large show comparable performance. At 60-70% of CG, FT-large SSL practically achieves the performance of LightGBM trained on 100% of CG. Further increase in CG size leads to a significant improvement in test AP for the pre-trained model. These observations suggest that SSL pre-training enables reducing the costs of fraud detection by either decreasing the CG size or increasing recall at the current precision level.
Also, it is worth noticing that the pre-trained FT-large significantly outperforms the one trained from scratch starting from 40-50% of CG. This indicates a good quality of representations learnt by the SSL methodology.
Figure 3. Test Average Precision for gradually increasing proportions of the Control Group. The horizontal dashed lines correspond to the large FT-Transformer (pre-trained with SSL) and LightGBM scores obtained on the full CG dataset.
@@ 6. Conclusions and Future Work
In this paper, we explored the recent advancements in tabular representation learning for the challenging task of fraud detection. In particular, motivated by the problem of selection bias, we scaled multiple pre-training strategies for the Transformer-powered deep learning model to large amounts of unlabeled data. According to various evaluation scenarios, the SSL pre-training strategy confidently outperforms one of the industry-standard GBDTs as well as Transformers trained from scratch in a supervised manner. Moreover, pre-training requires significantly less annotated data (up to 40%) to show a performance level of its GBDT competitor trained on the whole set of annotations.
@@ Acknowledgement
We would like to thank Maryam Moradbeigi for providing input and practical support to Bulat during his internship and Mozhdeh Ariannezhad for her contribution to the related work section.

### v2

@@ Abstract.
Transformer-based neural networks, empowered by Self-Supervised Learning (SSL), have demonstrated unprecedented performance across various domains. However, related literature suggests that tabular Transformers may struggle to outperform classical Machine Learning algorithms, such as Gradient Boosted Decision Trees (GBDT). In this paper, we aim to challenge GBDTs with tabular Transformers on a typical task faced in e-commerce, namely fraud detection. Our study is additionally motivated by the problem of selection bias, often occurring in real-life fraud detection systems. It is caused by the production system affecting which subset of traffic becomes labeled. This issue is typically addressed by sampling randomly a small part of the whole production data, referred to as a Control Group. This subset follows a target distribution of production data and therefore is usually preferred for training classification models with standard ML algorithms. Our methodology leverages the capabilities of Transformers to learn transferable representations using all available data by means of SSL, giving it an advantage over classical methods. Furthermore, we conduct large-scale experiments, pre-training tabular Transformers on vast amounts of data instances and fine-tuning them on smaller target datasets. The proposed approach outperforms heavily tuned GBDTs by a considerable margin of the Average Precision (AP) score in offline evaluations. Finally, we report the results of an online A/B experiment. Experimental results confirm the superiority of tabular Transformers compared to GBDTs in production, demonstrated by a statistically significant improvement in our business metric.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: conference details; Under Review ; isbn: XXXXX
@@ 1. Introduction
  • We investigate how pre-training affects the performance of FT-Transformer and its GBDT competitor in extensive offline experiments. We demonstrate that SSL pre-training consistently outperforms training a supervised model from scratch.
  • We report the results of an online A/B experiment comparing the performance of FT-Transformer and its GBDT competitor in production. We observe statistically significant improvements in terms of our key business metric.
@@ 2.1. Deep Learning for Tabular Data
Recently, multiple architectures extending Transformer self-attention to tabular data have been proposed aiming to challenge and eventually overcome Gradient Boosting-based methods (Yan et al., 2023; Kossen et al., 2021; Gorishniy et al., 2024; Hollmann et al., 2023; Song et al., 2019; Huang et al., 2020; Gorishniy et al., 2021). In 2019, Song et al. (Song et al., 2019) introduced the AutoInt architecture converting raw tabular features to embeddings that are passed through custom self-attention layers. TabTransformer (Huang et al., 2020), introduced in 2020, exploits conventional multi-head self-attention layers (Vaswani et al., 2017) for categorical feature embeddings, whereas numerical features are processed via a separate MLP branch of the network. The FT-Transformer architecture (Gorishniy et al., 2021) tokenizes each numerical and categorical feature, appends a class token, and applies the classical Transformer self-attention layers to the tokens. According to the original work, it outperforms the previously suggested Deep Learning models and even outperforms GBDTs on certain tasks.
@@ 4.1. Datasets
The biased (no-CG) dataset contains all instances from the traffic for which labels are available, except the ones sampled for the Control Group. The bias in this dataset is caused by a specific selection procedure, which introduces a shift in both features and labels, meaning that this dataset does not represent production data. The total number of instances in this dataset has the order of tens of millions.
The unbiased (CG) dataset, obtained using the Control Group strategy, contains a smaller amount of instances randomly sampled from the user traffic with reconstructed labels. The number of instances in this dataset is between one and two orders of magnitude less than the no-CG dataset size.
While the distribution of labels in the CG and no-CG datasets varies, the proportion of fraudulent examples in both datasets is very low. This results in highly imbalanced datasets. It is worth noting that we do not apply any data balancing technique. However, to derive the decision threshold we use an optimization algorithm that balances the costs and benefits of positive and negative samples. [^8] Because of confidentiality we cannot share more precise information.
@@ 5. Offline Evaluation
# Approach Backbone Pre-training Training Test AP
1 Sup. LightGBM - CG 0.471 $\pm$ 0.003
2 Sup. LightGBM - CG + no-CG 0.441 $\pm$ 0.013
3 Sup. + SR LightGBM All data CG + no-CG 0.442 $\pm$ 0.011
4 Sup. FT-base - CG 0.463 $\pm$ 0.013
5 DML FT-base CG + no-CG (+) CG 0.469 $\pm$ 0.017
6 SSL FT-base CG + no-CG (+) CG 0.469 $\pm$ 0.013
7 Sup. FT-large - CG 0.454 $\pm$ 0.009
8 DML FT-large CG + no-CG CG 0.488 $\pm$ 0.008
9 SSL FT-large CG + no-CG CG 0.491 $\pm$ 0.012
6. Online Evaluation
We conducted an online A/B experiment to further compare the performance of GBDTs with FT-Transformer for fraud detection at Booking.com. We followed the standard experiment protocol established within the scope of the production model. In the subsequent paragraphs, we describe the setup and the results of the experiment in as many details as possible, following internal company policies.
Evaluation metric. To evaluate online performance we use an internal utility-based business metric that aggregates positive and negative effects of decisions made by the model on each data instance, rewarding correct ones and penalizing wrong ones. The metric has a high variance, which we overcome through Trigger Analysis (Deng and Hu, 2015), i.e. tracking only the “triggered” part of the traffic where the base and the variant models make different decisions. In our case, Trigger Analysis yields a notably higher statistical power.
Base and variant models. The base model is a LightGBM trained and hyper-parameter tuned as described in Section 4.2, using SR weights (approach #3 in Table 2). This approach has been shown to outperform approach #1 in Table 2 in previous A/B experiments and has already been used in production. The variant model is the FT-Transformer trained with the Self-Supervised Learning strategy, as described in Section 4.3 (approach #9 in Table 2). Both models were trained and validated using the same data splits. We used the most recent data available at start of the of online experiment for training both models, which is similar to the datasets used in the offline experiments in terms of size and label distribution. Selecting the decision threshold was also performed using the same algorithm for both models (see Section 4.1).
Training resources. For training and hyperparameter tuning of the base model, we used a machine only with CPU resources. The variant model was trained on a machine with both GPU and CPU resources. For the same amount of data, and without performing any hyperparameter tuning for the variant, the training time is roughly six times higher for the variant.
A/B setup. Both variant and base models were deployed in the same way to our internal machine learning serving platform. They worked in production simultaneously for a period of time long enough to result in sufficient statistical power in the evaluation. To facilitate “Trigger Analysis”, both models were evaluated in production for each fraud detection instance, though the model deciding on the action was selected at random. It is worth noting that even though the real-time latency for the variant is around 20 times higher than for the base, it still falls under the budget we have for responding to live requests.
Results. The FT-Transformer variant outperforms the LightGBM baseline in terms of our business metric with a substantial margin. Specifically, we observe a 19.55% uplift in the triggered part of the traffic. The improvement is statistically significant, using an unpaired t-test with the significance level $\alpha=0.01$. It is worth mentioning that the additional training costs are negligible in light of improvements in the business metric.8
@@ 7. Conclusions and Future Work
In this paper, we explored the recent advancements in tabular representation learning for the challenging task of fraud detection. In particular, motivated by the problem of selection bias, we scaled multiple pre-training strategies for the Transformer-powered deep learning model to large amounts of unlabeled data. According to various offline evaluation scenarios, the SSL pre-training strategy confidently outperforms one of the industry-standard GBDTs as well as Transformers trained from scratch in a supervised manner. Moreover, we conducted online evaluation comparing FT-Transformer with LightGBM, our GBDT baseline, in an A/B experiment. We observed a statistically significant improvement over the baseline in terms of our internal utility-based business metric, which is inline with our offline findings.
@@ Acknowledgement
We would like to thank Maryam Moradbeigi for providing input and practical support to Bulat during his internship.
GenAI Usage Disclosure
Generative AI software tools were used to edit and improve the quality of the existing text in the same way we would use a typing assistant. They were also used in coding the same way we would use search engines and auto completion assistants.